OpenStudy (anonymous):

Find a number, b, such that f(g) = g(f) Where f(x) = 5x+b AND g(x) = 3x+4 Thanks in advance!

OpenStudy (anonymous):

5(3x+4)+b=3(5x+b)+4 15x+20+b=15x+3b+4 16=2b b=8

OpenStudy (anonymous):

Aha! Now it's coming back to me... thanks a lot!

OpenStudy (anonymous):

ur welcome :)